HI! how would i solve and equation like 487 / 6 * X +12-7=1628.3 find x?

Found 2 solutions by richwmiller, rothauserc:
Answer by richwmiller(17219)

Here is a great opportunity for a lesson in Mathematics and precision.
X and x are not the same in algebra.
There is no x in the given equation therefore is no solution.

I and i are not the same in English.

Besides that your presentation of the equation is not clear.
487 / 6 * X +12-7=1628.3 could mean several things
(487 / 6) *( X +12)-7=1628.3
487 /( 6 * X) +12-7=1628.3
You need to learn to write in the language of mathematics as well as in English.




Answer by rothauserc(4718)

487 / 6 * X +12-7=1628.3
order of operations do all multiplication and division in the order they appear from left to right, then addition and subtraction in the order they appear from left to right
81.2*x +5 = 1628.3
81.2x = 1623.3
x = 19.99
x = 20